The common stock of Dayton Repair sells for $43.19 a share. The stock is expected to pay $2.28 per share next year when the annual dividend is distributed. The firm has established a pattern of increasing its dividends by 2.15 percent annually and expects to continue doing so. What is the market rate of return on this stock? 


A. 7.59 percent  

B. 7.43 Percent  

C. 7.67 percent  

D. 7.14 percent  

E. 7.28 percent
Answer #1

rate of returns

= D1/P +G

D1= Dividend next year 2.28

P= price at present =43.19

G=growth rate of dividends =2.15%

= 2.28/43.19 +2.15%

=7.43%
Answer #2

Given :


P0 = 43.19 ($)

D1 = 2.28 ($)

g = 2.15 % = 0.0215



Now,


Dividend yield = D1/P0 

= 2.28/43.19

= 0.0528


Market rate of return, r 

= Dividend yield + g

= 0.0528+ 0.0215

= 0.0743

= 7.43 % 


So, market rate of return = 7.43 % (ANSWER).


answered by: Tulsiram Garg